OpenStudy (anonymous):

solve for x and y |x y| minus |-y x| |-y x| |x y | equals |4 6| |-4 6| does anyone know how to do this?

myininaya (myininaya):

ok so you have x+y=4, y-x=6, -y-x=-4, x-y=6 correct?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Its a matrix problem

myininaya (myininaya):

x+y=4 => y=4-x, then (4-x)-x=6 4-2x=6 => -2x=6/4=3/2 => x=-3/4 so y=4-x=4-(-3/4)=4+3/4=16/4+3/4=19/4

OpenStudy (anonymous):

im a little confused

myininaya (myininaya):

what bothers you?

OpenStudy (anonymous):

explain to me where the 2 came from? was it when you added the 2 ys and xs of the matrix?

myininaya (myininaya):

ok i took the first equation and solved it for y and then I wrote the second equation in terms of x so i only had one unknown in this equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

aw i see. thanks for your help!

myininaya (myininaya):

one sec okay?

myininaya (myininaya):

dont leave I have something for you

myininaya (myininaya):

I made a mistake. Click on this attachment,

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so x =-1. I didnt see what y equaled.

myininaya (myininaya):

you don't see y=4-(-1)=4+1=5

OpenStudy (anonymous):

ok i see it so y=5

myininaya (myininaya):

lol i ran out of room to put it on one line

OpenStudy (anonymous):

Thats ok i really appreciate your help. THANK YOU!

myininaya (myininaya):

np